OpenStudy (babystar16):

the sum of 2 numbers is 25. and their difference is 7. find the number.

OpenStudy (lukebluefive):

This is equivalent to the following series of equations: \[x + y = 25\] \[x = y - 7\] So, substitute the second equation into the first: \[(y - 7) + y = 25\] Simplify, and you'll have y. Then, plug that value of y into the second equation to get x.

OpenStudy (babystar16):

ty

OpenStudy (lukebluefive):

yw

OpenStudy (anonymous):

x+y=25 x-y=7....on solvng.. x=16,y=9

OpenStudy (babystar16):

ok got a new question

OpenStudy (babystar16):

the length of a rectangle is 5 more than twice the width. the permiter is 130 in. what is the area

OpenStudy (anonymous):

yea ask..

OpenStudy (anonymous):

5b+b=130 bredth=21.6 length=108.4

OpenStudy (anonymous):

area=lb=2341.4

OpenStudy (anonymous):

sory...b=10.8 i=54.2 area=587

OpenStudy (lukebluefive):

This equation gives the perimeter of a rectangle: \[2 \times length + 2 \times width\] In this case, the following is true as well: \[length = 2 \times width + 5\] Substitute this into the first equation and solve like the previous problem.

OpenStudy (lukebluefive):

Then multiply the length times the width to get the area, since: \[Area = length \times width\] Gives the area of a rectangle.

OpenStudy (anonymous):

hey u got the sol @babystar 16

OpenStudy (babystar16):

ty